OpenStudy (anonymous):

how do i find n 2/3+n+6=3/4

OpenStudy (anonymous):

2/3+n+6=3/4 ===> n= (3/4)-(2/3) -6 ===> n =-11/12

OpenStudy (anonymous):

write it with equation forum please Santi890

OpenStudy (anonymous):

\[n=-\frac{71}{12} \]

OpenStudy (anonymous):

raheen got the work right but the fraction arithmetic wrong. robtobey got the right answer.